prettygirl15:

2x^2+8x+8=0 A) 0; two imaginary solutions B) 0; one real solution C) 65; two real solutions D) 0; two real solutions

SpiderGwen:

Two real solutions.

itzlanalol:

@spidergwen wrote:
Two real solutions.
your not supposed to give a full answer. you have to work through it with them

SpiderGwen:

@itzlanalol wrote:
@spidergwen wrote:
Two real solutions.
your not supposed to give a full answer. you have to work through it with them
Oops.

SpiderGwen:

🙄

itzlanalol:

@spidergwen wrote:
🙄
and your wrong btw

SpiderGwen:

@itzlanalol wrote:
@spidergwen wrote:
🙄
and your wrong btw
I found two real solutions.

itzlanalol:

@spidergwen wrote:
@itzlanalol wrote:
@spidergwen wrote:
🙄
and your wrong btw
I found two real solutions.
it's only one solution

xxDeppressionxx:

Please refrain from giving direct answers that it strictly against the rules, please delete your comments and explain how you got the answer. Thank you🙏.

itzlanalol:

STEP 1: (0 - 2x2) - 8x) - 8 = 0 STEP 2: -2x2 - 8x - 8 = -2 • (x2 + 4x + 4) step 3: -2 • (x + 2)2 = 0 One solution was found : X=-2

xxDeppressionxx:

@itzlanalol wrote:
STEP 1: (0 - 2x2) - 8x) - 8 = 0 STEP 2: -2x2 - 8x - 8 = -2 • (x2 + 4x + 4) step 3: -2 • (x + 2)2 = 0 One solution was found : X=-2
Thank you🙏.

itzlanalol:

@xxdeppressionxx wrote:
@itzlanalol wrote:
STEP 1: (0 - 2x2) - 8x) - 8 = 0 STEP 2: -2x2 - 8x - 8 = -2 • (x2 + 4x + 4) step 3: -2 • (x + 2)2 = 0 One solution was found : X=-2
Thank you🙏.
your welcome

xxDeppressionxx:

Can you also site you source.

xxDeppressionxx:

Of where you got your answer from.

itzlanalol:

of course yes

itzlanalol:

It explains everything

prettygirl15:

@itzlanalol wrote:
STEP 1: (0 - 2x2) - 8x) - 8 = 0 STEP 2: -2x2 - 8x - 8 = -2 • (x2 + 4x + 4) step 3: -2 • (x + 2)2 = 0 One solution was found : X=-2
2 isn't a oppstion on there

xxDeppressionxx:

The answer as they were saying is one solution

xxDeppressionxx:

@itzlanalol wrote:
STEP 1: (0 - 2x2) - 8x) - 8 = 0 STEP 2: -2x2 - 8x - 8 = -2 • (x2 + 4x + 4) step 3: -2 • (x + 2)2 = 0 One solution was found : X=-2

prettygirl15:

A) 0; two imaginary solutions B) 0; one real solution C) 65; two real solutions D) 0; two real solutions these are the only opposition

xxDeppressionxx:

What do you think one solution means?

xxDeppressionxx:

One solution means one real solution in math terms because its a real life math problem.

itzlanalol:

@prettygirl15 wrote:
A) 0; two imaginary solutions B) 0; one real solution C) 65; two real solutions D) 0; two real solutions these are the only opposition
it means that it'd be one solution with the reasons we both gave you

jhonyy9:

so to solve this math problem easy right just you need calcule the discriminante and use these 3 cases D < 0 D = 0 D > 0 and you need to know how are these solutions (roots) of this quadratic in these 3 cases of discriminant